OpenStudy (anonymous):

The sum of four integers is 287 . The three smaller are consecutive integers and the two larger are consecutive even integers. Find the product of the smallest and the largest of the four integers.

OpenStudy (anonymous):

What have you attempted to solve your problem?

OpenStudy (anonymous):

I got 5180 for an answer I want someone to check.

OpenStudy (anonymous):

Okay then, just a second.

OpenStudy (anonymous):

Let x be the smallest one. Hence the the others are: $$x + 1,~x + 2,~x +4$$ Therefore one has to solve the equation $$x + x + 1 + x + 2 + x +4 = 287$$ $$4x + 7 = 287$$ $$4x = 287 - 7 = 280$$ $$x = 280/4 = 70$$ \(\textbf{Hence the answer is:}\) $$70 \times (70 + 4) = 5180$$ You are correct.

OpenStudy (anonymous):

Awesome thank you!